a jet flew from tokyo to bangkok, a distance of 4800km. on return trip, the speed was decreased by 200 km/h. if the difference in times f the flights was 2 hours, what was the jet's speed from bangkok to tokyo?
Answer by richwmiller(17219)

r*t=4800,
(r-200)*(t+2)=4800
negative answers are rejected
r = 800, t = 6
The jet went at 600 kph from bangkok to tokyo for 8 hours